OpenStudy (anonymous):

Which is 6^3 × 6^4 in exponential form? 6^7? @Hero @SolomonZelman

OpenStudy (anonymous):

@Hero @SolomonZelman @whpalmer4 @mathstudent55 @satellite73 @Yuba

OpenStudy (anonymous):

@Loser66 @SolomonZelman @matricked @mathslover @mathlover2014 @mathstudent55 Am I right? Is it 6^7?

OpenStudy (solomonzelman):

\(\Large\color{purple}{ \bf a^b \times a^c=a^{b+c} }\) \(\Large\color{purple}{ \bf a^b \div a^c=a^{b-c} }\) \(\Large\color{purple}{ \bf 3^2 \times 3^5=3^{2+5} =3^7=2,187 }\) \(\Large\color{purple}{ \bf 7^6 \times 7^2=7^{6+2} =7^8=5,764,801 }\) \(\Large\color{purple}{ \bf 4^4 \div 4^2=4^{4-2} =4^2=16 }\) \(\Large\color{purple}{ \bf 6^4 \div 6^3=6^{3-2} =6^1=6 }\)

OpenStudy (solomonzelman):

Hope this is helpful :)

OpenStudy (anonymous):

A. 36^12 B. 7^6 C. 612 D. 6^7 Er....I just want to know if I was right. :) @SolomonZelman

OpenStudy (anonymous):

@mathslover Was I right? :):):)

OpenStudy (solomonzelman):

Yes !

OpenStudy (solomonzelman):

Have a good night, going offline in 4.5 minutes.